OpenStudy (anonymous):

find the solutions to the systems of nonlinear equations: y=2x x^2+y^2=15

OpenStudy (anonymous):

substitute and solve

OpenStudy (anonymous):

y = 2x => x^2 +(2x)^2 = 15

OpenStudy (anonymous):

I think I am doing that but I dont know if I'm getting the right answer or working the problem right

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

x = square root of 3 and negative square root of 3 y = 2 square root of 3 and -2 square root 3

OpenStudy (anonymous):

there you go... except it should be two points so write them as pairs (x, y)

OpenStudy (anonymous):

Oh okay sweet :D thanks for the tip ! could you help me with one more problem similar to this one?

OpenStudy (anonymous):

let's see

OpenStudy (anonymous):

x^2 + y^2= 25 x^2 -y^2 --- --- = 1 5^2 5^2 I don't know on which equation to start with

OpenStudy (anonymous):

solve the first for one of the variables and plug it into the second.

OpenStudy (anonymous):

So move the y^2 to the other side then square root everything?

OpenStudy (anonymous):

no need to sqrt.

OpenStudy (anonymous):

so just flip the y^2 to the other side? x^2 = 25 - y^2

OpenStudy (anonymous):

yeah, that'll work

OpenStudy (anonymous):

so replace the x^2 with 25 - y^2

OpenStudy (anonymous):

you can do that... i'd do y^2 = 25 - x^2

OpenStudy (anonymous):

how come is it easier to do it with y first?

OpenStudy (anonymous):

it really doesn't matter

OpenStudy (anonymous):

oh ok I thought you HAD to start with y first lol

OpenStudy (anonymous):

no

OpenStudy (anonymous):

going for dinner, i'll be back later

OpenStudy (anonymous):

oh ok im working through the problem right now

OpenStudy (anonymous):

you should get 4 points

OpenStudy (anonymous):

you multiply the denominator by 25 right? or 5

OpenStudy (anonymous):

25